# Conceptualization, Data Curation

### v2

@@ Abstract
Condition monitoring plays a significant role in the safety and reliability of modern industrial systems. Artificial intelligence (AI) approaches are gaining attention from academia and industry as a growing subject in industrial applications and as a powerful way of identifying faults. This paper provides an overview of intelligent condition monitoring and fault detection and diagnosis methods for industrial plants with a focus on the open-source benchmark Tennessee Eastman Process (TEP). In this survey, the most popular and state-of-the-art deep learning (DL) and machine learning (ML) algorithms for industrial plant condition monitoring, fault detection, and diagnosis are summarized and the advantages and disadvantages of each algorithm are studied. Challenges like imbalanced data, unlabelled samples and how deep learning models can handle them are also covered. Finally, a comparison of the accuracies and specifications of different algorithms utilizing the Tennessee Eastman Process (TEP) is conducted. This research will be beneficial for both researchers who are new to the field and experts, as it covers the literature on condition monitoring and state-of-the-art methods alongside the challenges and possible solutions to them.
@@
{highlights}
A comprehensive review of intelligent condition monitoring methodologies for industrial plants is conducted.
Different deep learning and neural network-based approaches are fairly compared on a standard benchmark dataset.
Challenges regarding the data quality, explainability and possible solutions for tackling them are explored.
@@ 1 Introduction
Condition Monitoring (CM) is crucial for any industry as it improves the safety, reliability and efficiency of the systems [80]. In industrial systems, CM can be defined as the process of analyzing measurable variables to identify significant changes which are signs of a developing fault. A fault is an unpermitted deviation of the characteristic property of a variable from standard behaviour [34]. Faults can be categorized into three classes based on their time dependency, abrupt fault (stepwise), incipient fault (drift-like), and intermittent fault (happens at usually irregular intervals) shown in Figure 1. If the fault is not detected and corrected on time, it may lead to a failure which is a permanent interruption in the system performance. The tasks of fault detection and diagnosis (FDD) have various definitions in different literature; however, in this paper, we define these tasks as fault detection, fault classification, fault isolation, and fault identification, based on [34] and [35]. Fault detection is the task of determining if a fault has occurred in the system or not. Fault classification determines the type of faults, and discriminates different groups of faults. Fault diagnosis identifies the type of faults with as much specificity as possible, including the problem size, location, and time of detection. Fault isolation determines the location of each fault, for example, which part of the system has become faulty. Fault identification determines the type, magnitude and cause of the fault. In general FDD methods can be divided into three main categories, hardware redundancy, model-based fault detection, and signal-based fault detection. Hardware redundancy is a method in which multiple sensors, hardware, and software are used to measure variables. With these measurements, final FDD decisions can be made based on a weighted vote among the collected information. This method is highly reliable and common in very sensitive systems, however, it requires additional space to place the redundant hardware and can also be expensive. In model-based fault detection, faults are detected by comparing the available measurements of the system in conjunction with information from a mathematical model of the system. A residual amount is determined using the difference between actual measurements and estimates obtained from the model. Then, by applying a threshold to this residual quantity, a fault can be detected. Although this method is reliable, defining a precise model of a complex system is not an easy task, posing a challenge even to domain experts. On the other hand, signal-based methods rely solely on the analysis of data collected from sensors to detect deviations from normal behaviour and are the most frequently employed FDD methods. This approach is useful for monitoring complex systems where traditional approaches may not be sufficient, and it has been widely applied to many different industrial systems and processes. As there are numerous sensors and data in any industrial system, deep learning algorithms can be utilized to handle such information. However, this approach is highly dependent on the quality and availability of the data. The focus of this paper is on studying different state-of-the-art deep learning algorithms that have been used for the condition monitoring of industrial plants with a focus on chemical processing plants. A comparison between different approaches is established through the analysis of their application on the Tennessee Eastman Process (TEP) simulator benchmark. This paper covers different deep learning methods for condition monitoring of industrial plants and compares and summarizes these methods, also discusses the common problems of data-driven condition monitoring and introduces solutions to address them.
Figure 1: Time-depandancy of faults [34].
Condition monitoring is an inevitable task in industrial chemical systems and as such, numerous studies have been conducted in this topic. Yin et al. [108] compared basic data-driven fault diagnosis and process monitoring methods such as Principal component analysis (PCA), DPCA, two variants of Partial least squares (PLS), Fisher discriminant analysis (FDA) and Independent component analysis (ICA) on the TEP benchmark. Their study compares both the performance and the computational complexity of the examined methods. Another study by Md Nor et al. [55] provides a comprehensive review of data-driven condition monitoring methods in chemical process systems. The research covers multivariate statistical analysis such as PCA, PLS, ICA, and FDA. In addition, the work also examines machine learning approaches such as artificial neural networks (ANNs), neuro-fuzzy (NF) methods, support vector machines (SVMs), Gaussian mixture models (GMMs), K-nearest neighbours algorithm (KNN), and Bayesian networks (BNs).
Limited studies have reviewed data-driven and machine-learning FDD methods for complex industrial processes[108, 55, 109, 8]. However, to the best of our knowledge, there is no comprehensive study on deep learning and neural network-based CM approaches that fairly compares the methods based on a common benchmark case study. In this research, intelligent CM algorithms have been thoroughly analyzed and fairly compared within the context of industrial chemical processes, covering implementation challenges such as the explainability of AI methods and data scarcity. Additionally, this work considers the current state and future directions of research within the broader scope of industrial CM.
1.1 Performance Metrics
@@
Performance and evaluation metrics serve as valuable tools for comparing and evaluating fault detection and diagnosis capabilities of different methods. This section describes the most important and widely used metrics within the context of industrial CM to provide sufficient knowledge for comparing such studies throughout the remainder of this paper. Many of these studies include the classification terms True Positive (TP), False Positive (FP), True Negative (TN), and False Negative (FN). TP represents the samples that belong to the normal class and are correctly classified as normal, FP denotes misclassifications of faulty samples as normal, TN represents the faulty samples that are correctly classified as faulty and FN refers to the incorrect classification of normal samples as faulty. It is crucial to achieve accurate detection of both normal and faulty samples. However, minimising false positives can be considered a more important task since it shows that no fault condition is mistakenly considered normal. If there is any abnormality, this prioritisation enables prompt corrective action.
Accuracy quantifies the number of correctly classified data samples out of the total data sample count, as expressed in Equation 1. It is important to note that accuracy may not be the most suitable metric when dealing with imbalanced datasets.
$\begin{split}Accuracy=\frac{TP+TN}{TP+TN+FP+FN}\end{split}$ (1)
Precision evaluates the quality of a positive prediction made by the model and is defined by Equation 2.
$\begin{split}Precision=\frac{TP}{TP+FP}\end{split}$ (2)
Recall also known as sensitivity or true positive rate measures the number of the predicted positive instances out of all positive examples within the considered data and is expressed in Equation 3.
$\begin{split}Recall=\frac{TP}{TP+FN}\end{split}$ (3)
Fault detection rate (FDR), reflects the proportion of actual faults that are successfully detected by the system and is defined in equation 4.
$\begin{split}FDR=\frac{TN}{FP+TN}\end{split}$ (4)
False alarm rate (FAR), also known as the false positive rate indicates the proportion of falsely predicted positive instances and is presented in equation 5.
$\begin{split}FAR=\frac{FP}{FP+TN}\end{split}$ (5)
The F1-score as defined in equation 6 combines both precision and recall to provide a more balanced metric for model evaluation.
$\begin{split}F1_{-}Score=2\times\frac{Precision\times Recall}{Precision+Recall}\end{split}$ (6)
@@ 2 Research Methodology
This paper focuses on signal-based approaches for condition monitoring of industrial plants. As the amount of available sensor data can be abundant in real-world scenarios, the main focus of this research is on machine learning and deep learning techniques, especially neural networked-based approaches that can handle large amounts of data. For a fair comparison of the different applied methods, a simulated plant known as the Tennessee Eastman Process (TEP) is considered as the case study and different condition monitoring methods are compared based on their performance over this use case. A three-level keyword assembly is used to gather related papers and research to address the mentioned aspects. Table 1 shows the keyword breakdown of the query. All papers containing "Condition Monitoring", "Fault Detection", "Fault Diagnosis", "Prediction", "Fault Prognosis", and "Anomaly Detection" can be considered for review as they cover the application aspect of this paper. From the methodological point of view, the "Machine Learning" and "Deep Learning" approaches with a focus on "Neural Networks" have been selected. More than 400 papers are extracted from the Web of Science database using this query. The number of publications over the years is shown in Figure 2. Early papers on deep learning in this field were published in the early 2000s; however, these methods gained significant attention in recent years, especially after 2015, when research on more advanced machine learning and deep learning algorithms increased in popularity. This increase in popularity is largely attributed to improved computing resources, algorithmic breakthroughs, and the ability of such methods to handle complex big data. Within the scope of ML and DL, many different algorithms have been used for condition monitoring of industrial plants. Of the selected papers, the most frequently used algorithms are Autoencoders, Convolutional Neural Networks (CNNs), Recursive Neural Networks (RNNs), Deep Beleif Networks (DBNs), Attention mechanisms and Generative Adversarial Networks (GANs). Figure 3 shows the distribution of different algorithms among the studied papers.
Figure 2: Number for publications over years.
Figure 3: Distribution of Methods.
Table 1: The proposed keywords and research queries
Context Query Search Fields
@@
Application "Condition Monitoring" or "Fault detection" or "Fault diagnosis" or "Prediction" or "Fault Prognosis" or "Anomaly Detection" Title & Keywords & Abstract
Method "Machine Learning" or "Deep Learning" or "Neural Network" Title & Keywords & Abstract
Case Study "Tennessee Eastman" Title & Keywords & Abstract
3 Tennessee Eastman Process
@@
The Tennessee Eastman Process (TEP) is a realistic simulation of an industrial chemical plant which is the most used benchmark for process control and monitoring studies [56, 108]. The process is described in [22], and both the Fortran and Matlab simulation codes, along with numerous generated datasets, can be found online. The process consists of five main units: the reactor, the product condenser, a vapour-liquid separator, a recycle compressor and a product stripper. The flow diagram of the process is shown in Figure 4. The process produces two products from four reactants, along with an inert and a byproduct. This results in a total of eight components denoted as A, B, C, D, E, F, G, and H. The reactions are shown as follows:
@@
The original simulation consists of 41 process variables and 12 manipulated variables, with sampling intervals of 3, 6 and 15 minutes. A total of 21 process faults, including step, random variation, slow drift and sticking faults, are defined for this simulation [108]. Refer to Table 2 for details. Some faults like 3, 9 and 15 are harder to detect.
Figure 4: Tennessee Eastman Process flow diagram adapted from [22]
@@
IDV(16) Unknown Unknown
IDV(17) Unknown Unknown
IDV(18) Unknown Unknown
IDV(19) Unknown Unknown
IDV(20) Unknown Unknown
@@
4 Intelligent Condition Monitoring Approaches
This section explores some of the most important algorithms that have been successfully used for the condition monitoring of industrial plants, with a focus on the TEP.
4.1 Recurrent Neural Network (RNN) and Long Short-Term Memory (LSTM)
Recurrent Neural Networks (RNN), are the networks that can learn time dependencies between data samples. These networks have feedback connections, enabling them to capture information from previous steps. The model is usually trained with back-propagation and suffers from gradient vanishing/exploding. Long Short-Term Memory (LSTM) networks are improved RNNs, consisting of memory cells and the gating mechanism for capturing information over a long time. The forget gate handles the gradient vanishing problem [31]. Gated recurrent unit (GRU) solves the vanishing gradient problem like LSTM but with fewer gates and no separate internal memory. RNN, LSTM and GRU structures are shown in Figure 5.
Figure 5: RNN vs LSTM vs GRU.[81]

### v3

# Visualization, Writing - Review

@@ Abstract
Condition monitoring is essential for ensuring the safety, reliability, and efficiency of modern industrial systems. With the increasing complexity of industrial processes, artificial intelligence (AI) has emerged as a powerful tool for fault detection and diagnosis, attracting growing interest from both academia and industry. This paper provides a comprehensive overview of intelligent condition monitoring methods, with a particular emphasis on chemical plants and the widely used Tennessee Eastman Process (TEP) benchmark. State-of-the-art machine learning (ML) and deep learning (DL) algorithms are reviewed, highlighting their strengths, limitations, and applicability to industrial fault detection and diagnosis. Special attention is given to key challenges, including imbalanced and unlabeled data, and to strategies by which models can address these issues. Furthermore, comparative analyses of algorithm performance are presented to guide method selection in practical scenarios. This survey is intended to benefit both newcomers and experienced researchers by consolidating fundamental concepts, summarizing recent advances, and outlining open challenges and promising directions for intelligent condition monitoring in industrial plants.
@@ 1 Introduction
Condition monitoring (CM) is fundamental to industrial systems, enabling early detection of faults and supporting efficient, safe, and reliable operations Shojaeinasab et al. (2022). Within this context, fault detection and diagnosis (FDD) reinforces process safety by enabling the timely isolation of system faults, helping to prevent escalation into incidents that could affect human safety or the environment. In industrial systems, CM can be defined as the process of analyzing measurable variables to identify significant changes that are signs of a developing fault. A fault is an unpermitted deviation of the characteristic property of a variable from standard behavior Isermann (2005). Faults can be categorized into three classes based on their time dependency: abrupt fault (stepwise), incipient fault (drift-like), and intermittent fault (usually occurs at irregular intervals), shown in Figure 1. If the fault is not detected and corrected on time, it may lead to a failure, which is a permanent interruption in the system’s performance. The tasks of FDD have various definitions in different literature; however, in this paper, we define these tasks as fault detection, fault classification, fault isolation, and fault identification, based on Isermann (2005) and Izadi Najaf Abadi (2006).
@@
Fault detection is the task of determining whether a fault has occurred in the system or not. Fault classification determines the type of faults and discriminates between different groups of faults. Fault diagnosis identifies the type of faults with as much specificity as possible, including problem size, location, and time of detection. Fault isolation determines the location of each fault, for example, which part of the system has become faulty. Fault identification determines the type, magnitude, and cause of the fault. In general, FDD methods can be divided into three main categories: hardware redundancy, model-based fault detection, and signal-based fault detection. Hardware redundancy is a method in which multiple sensors, hardware, and software are used to measure variables. With these measurements, final FDD decisions can be made based on a weighted vote among the collected information. This method is highly reliable and common in very sensitive systems; however, it requires additional space to place the redundant hardware and can also be expensive. In model-based fault detection, faults are detected by comparing the available measurements of the system in conjunction with information from a mathematical model of the system. A residual amount is determined using the difference between the actual measurements and the estimates obtained from the model. Then, by applying a threshold to this residual quantity, a fault can be detected. Although this method is reliable, defining a precise model of a complex system is not an easy task, posing a challenge even to domain experts. On the other hand, signal-based methods rely solely on the analysis of data collected from sensors to detect deviations from normal behavior and are the most frequently employed FDD methods. This approach is useful for monitoring complex systems where traditional approaches may not be sufficient, and it has been widely applied to many different industrial systems and processes. As there are numerous sensors and data in any industrial system, deep learning algorithms can be utilized to handle such information. However, this approach is highly dependent on the quality and availability of the data. ML algorithms are widely used for condition monitoring Surucu et al. (2023). Moreover, combining the mentioned FDD methods, using hybrid approaches, and integrating prior knowledge into the data-driven and ML methods will enhance the accuracy and safety of the process.
Figure 1: Time-dependency of faults Isermann (2005).
@@
Performance and evaluation metrics serve as valuable tools for comparing and evaluating fault detection and diagnosis capabilities of different methods. This section describes the most important and widely used metrics within the context of industrial CM to provide sufficient knowledge for comparing such studies throughout the remainder of this paper. Many of these studies include the classification terms True Positive (TP), False Positive (FP), True Negative (TN), and False Negative (FN). In this study, we treat the positive class as fault / abnormal and the negative class as normal:
  • TP (true positives): samples that are actually faulty and predicted as faulty.
  • TN (true negatives): samples that are actually normal and predicted as normal.
  • FP (false positives): samples that are actually normal but predicted as faulty (false alarms).
  • FN (false negatives): samples that are actually faulty but predicted as normal (missed detections).
It is crucial to achieve accurate detection of both normal and faulty samples. However, minimizing false negatives is considered a more important task, as it ensures that no fault condition is mistakenly considered normal. If there is any abnormality, this prioritization enables prompt corrective action. The most used performance metrics are shown in Table 1.
Table 1: Common performance metrics for fault detection and diagnosis
Metric Formula Description and Notes
Accuracy $\frac{TP+TN}{TP+TN+FP+FN}$ Measures the overall correctness of the model.Proportion of correctly classified samples out of all samples. It may be misleading for highly imbalanced datasets.
Precision $\frac{TP}{TP+FP}$ Indicates how many of the predicted faulty conditions were actually faulty. Higher precision means fewer false alarms.
Recall (Sensitivity) $\frac{TP}{TP+FN}$ Measures the model’s ability to find all actual faulty conditions (true positive rate). Useful in safety-critical systems.
False Discovery Rate (FDR) $\frac{FP}{TP+FP}$ The proportion of false alarms among all positive predictions. It’s the complement of precision. A low FDR is desirable as it indicates a lower rate of erroneous fault detections.
False Alarm Rate (FAR) $\frac{FP}{FP+TN}$ The proportion of actual healthy conditions that were incorrectly identified as faulty. Lower values indicate better reliability (also known as false positive rate).
F1-score $2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ Harmonic mean of precision and recall. Recommended for imbalanced datasets.
The focus of this paper is on studying different state-of-the-art deep learning algorithms that have been used for the condition monitoring of industrial plants with a focus on chemical processing plants. A comparison between different approaches is established through the analysis of their application on the Tennessee Eastman Process (TEP) simulator benchmark. This paper covers different deep learning methods for condition monitoring of industrial plants, compares and summarizes these methods, discusses the common problems of data-driven condition monitoring, and introduces solutions to address them.
Condition monitoring is an inevitable task in industrial chemical systems, and as such, numerous studies have been conducted on this topic. Yin et al. Yin et al. (2012) compared basic data-driven fault diagnosis and process monitoring methods such as Principal component analysis (PCA), DPCA, two variants of Partial least squares (PLS), Fisher discriminant analysis (FDA), and Independent component analysis (ICA) on the TEP benchmark. Their study compares both the performance and the computational complexity of the examined methods. Another study by Md Nor et al. Md Nor et al. (2020) provides a comprehensive review of data-driven condition monitoring methods in chemical process systems. The research covers multivariate statistical analysis such as PCA, PLS, ICA, and FDA. In addition, the work also examines machine learning approaches such as artificial neural networks (ANNs), neuro-fuzzy (NF) methods, support vector machines (SVMs), Gaussian mixture models (GMMs), K-nearest neighbours algorithm (KNN), and Bayesian networks (BNs).
Limited studies have reviewed data-driven and machine-learning FDD methods for complex industrial processes Yin et al. (2012); Md Nor et al. (2020); Yin et al. (2014); Bi et al. (2022). However, to the best of our knowledge, there is no comprehensive study on deep learning and neural network-based CM approaches that fairly compares the methods based on a common benchmark case study. In this research, intelligent CM algorithms have been thoroughly analyzed and compared within the context of industrial chemical processes, covering implementation challenges such as the explainability of AI methods and data scarcity. Common challenges in data-driven condition monitoring of industrial systems, such as unlabeled, unseen, and imbalanced data, and how to tackle them, are introduced. Solutions for dealing with different types of data limitations and the explainability of ML methods are also explored. Additionally, this work considers the current state and future directions of research within the broader scope of industrial CM. This research offers value to both new researchers and experts in the field. It introduces fundamental concepts for those new to intelligent condition monitoring while providing a comprehensive comparison of recent advancements and a discussion of challenges and limitations to guide future research.
@@ 2 Research Methodology
This study focuses on signal-based approaches for condition monitoring of industrial plants. , where the availability of rich sensor data in real-world applications motivates the use of machine learning (ML) and deep learning (DL) techniques. In particular, neural network–based methods have attracted increasing attention due to their ability to process high-dimensional data and capture complex nonlinear relationships. To provide a consistent basis for comparison, the Tennessee Eastman Process (TEP), a widely used benchmark in process systems engineering, is selected as the primary case study. Performance of different condition monitoring methods is reviewed in the context of this benchmark, highlighting both their strengths and limitations. A systematic literature search was conducted using a three-level keyword assembly designed to capture research on intelligent condition monitoring, AI techniques, and benchmark applications. The search query using All Fields operators was formulated as follows:
@@
("Condition Monitoring" OR "Fault Detection" OR "Fault Diagnosis" OR "Prediction", "Fault Prognosis" OR "Anomaly Detection") AND ("Machine Learning" OR "Deep Learning" OR "Neural Networks") AND ("Tennessee Eastman").
Using this formulation, more than 450 papers were retrieved from the Web of Science (WoS) database. The inclusion criteria required papers to be written in English, focused on ML/DL-based signal-level condition monitoring of industrial plants. Exclusion criteria included papers that focused solely on traditional non-neural ML methods (e.g., PCA, SVM) without DL integration, or lacked sufficient methodological detail for meaningful comparison. The temporal distribution of the selected papers is shown in Figure 2(a). While early applications of ML to condition monitoring emerged in the early 2000s, interest in advanced DL methods accelerated after 2015, largely due to algorithmic breakthroughs, improved computational resources, and the availability of large-scale industrial datasets. Within the scope of ML and DL, the most frequently employed algorithms include Autoencoders (AEs), Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Deep Belief Networks (DBNs), attention mechanisms, and Generative Adversarial Networks (GANs). Their distribution across the exdtracted papers is illustrated in Figure 2(b). These categories form the backbone of the comparative analysis presented in this review.
(a) Number of publications over the years.
(b) Visual taxonomy of methods.
Figure 2: Review metrics.
3 The Evolution of Intelligent Condition Monitoring
This section explores some of the most important algorithms that have been successfully used for the condition monitoring of industrial plants, with a particular focus on data-driven methods. These approaches leverage machine learning techniques to detect and diagnose faults directly from process signals, often without relying on detailed physical models. While the emphasis is on recent advances, we begin with a brief overview of foundational methodologies, including probabilistic reasoning and hybrid diagnostic systems, which have shaped the development of modern intelligent monitoring frameworks.
3.1 Early Methodologies and Foundations
Early work in condition monitoring laid the foundation for modern data-driven fault detection strategies through the use of interpretable, statistically grounded models. Among the most influential were multivariate statistical process monitoring techniques such as PCA and PLS, which enabled unsupervised and supervised monitoring of high-dimensional process data Kresta et al. (1991); Wilson and Irwin (2000). Extensions like Dynamic PCA Ku et al. (1995) and Multiway PCA Nomikos and MacGregor (1994) introduced temporal structure, improving detection sensitivity for dynamic and batch processes.
As limitations of linear models became evident, researchers explored nonlinear extensions and hybrid methods. For example, Nonlinear Dynamic PCA combined neural networks with PCA to model nonlinear dynamics Lin et al. (2000), while Canonical Variate Analysis (CVA) offered improved fault sensitivity in time-dependent settings Russell et al. (2000). Hybrid approaches, such as the combination of signed directed graphs with dynamic PLS, allowed for structured fault isolation and improved diagnosis of multiple simultaneous faults Lee et al. (2004). These methods emphasized the value of integrating statistical learning with process-specific knowledge.
Probabilistic and knowledge-informed techniques also emerged during this period. Early applications of Bayesian reasoning, expert systems, and qualitative models provided interpretable and explainable fault analysis Venkatasubramanian et al. (2003). Notably, Kulkarni et al. demonstrated that embedding domain insights into support vector machines improved classification accuracy, suggesting a transition toward learning approaches informed by prior knowledge Kulkarni et al. (2005).
The ideas introduced by these early approaches continue to influence the development of more advanced monitoring frameworks. In the following sections, we examine how recent advances build upon and extend these early contributions to address increasingly complex monitoring challenges.
3.2 Prominent Process Monitoring Benchmarks
Several simulators and datasets serve as benchmarks for process condition monitoring, including the Tennessee Eastman Process (TEP), PenSim, RAYMOND, DAMADICS, IndPenSim, FCC fractionator, Cranfield Multiphase Flow Facility Melo et al. (2022). In this research, TEP is utilized as the primary benchmark for comparison due to its widespread acceptance and use in numerous studies. Some of the other case studies are examined in the section titled Industrial Applicability Beyond TEP.
@@ 3.2 Prominent Process Monitoring Benchmarks
TEP is a realistic simulation of an industrial chemical plant, which is the most used benchmark for process control and monitoring studies Melo et al. (2022); Yin et al. (2012). The process is described in Downs and Vogel (1993), and both the Fortran and Matlab simulation codes, along with numerous generated datasets, can be found online. The process consists of five main units: the reactor, the product condenser, a vapour-liquid separator, a recycle compressor, and a product stripper. The flow diagram of the process is shown in Figure 3. The process produces two products from four reactants, along with an inert and a byproduct. This results in a total of eight components denoted as A, B, C, D, E, F, G, and H. The reactions are shown as follows:
@@
The original simulation consists of 41 process variables and 12 manipulated variables, with sampling intervals of 3, 6, and 15 minutes. A total of 21 process faults, including step, random variation, slow drift, and sticking faults, are defined for this simulation Yin et al. (2012). Refer to Table 2 for details. Some faults, like 3, 9, and 15, are harder to detect.
Figure 3: Tennessee Eastman Process flow diagram adapted from Downs and Vogel (1993)
@@
IDV(16 -20) Unknown Unknown
@@
3.3 Recurrent Neural Networks
Recurrent neural networks (RNNs) are a class of deep learning models specifically designed to capture temporal dependencies in sequential data, making them well-suited for condition monitoring in dynamic industrial processes. Unlike feedforward networks, RNNs contain feedback connections that allow information to persist across time steps. However, standard RNNs often suffer from issues such as vanishing or exploding gradients during training. To address these limitations, Long Short-Term Memory (LSTM) networks were introduced with gated memory cells that enable learning over longer time horizons Hochreiter and Schmidhuber (1997). Gated Recurrent Units (GRUs) offer a more lightweight alternative to LSTMs by simplifying the gating mechanisms, often achieving comparable performance with reduced computational overhead.